LyssaM:

On Monday, a company stock closed at a price of $30.80 per share. On Tuesday, it increased by $1.20 per share. On Wednesday, it fell by 6.25%. On Thursday, it fell by another 4%, and on Friday, it increased by 2.5%. What is the net change in the stock price from Monday to Friday?

LyssaM:

I don't understand this

justjm:

Go step by step. On Monday, it was $30.80 per share. net = $30.80 Tuesday, it increased by 1.20 per share. net = 30.80 + 1.20 On Wednesday, it decreased by 6.25 percent. net = 0.9375(30.80 + 1.20) On thurs, it fell by another 4 percent. net = 0.96(0.9375(30.80+1.20)) On friday, it increased by 2.5%. net = 1.025(0.96(0.9375(30.80+1.20))) So by the time friday comes, the final price is something like this: \((1.025(0.96(0.9375(30.80+1.20))) \) Calculate that. But that isn't the answer, it is asking for the net change. So substract the value of Monday from Friday